OpenStudy (anonymous):

A block initially at rest is pushed horizontally along a frictionless surface with a constant force. After moving a distance d the speed of the block is v. What is the speed of the block after being pushed a distance 2d?

OpenStudy (anonymous):

Find the acceleration from: V2^2 = V1^2 + 2ad with V1 = 0 Then you can find the velocity at 2d from the same equation.

OpenStudy (anonymous):

v=sqrt(2 a d) thus after 2d you get sqrt(2)*v